OpenStudy (anonymous):

How do I solve x + y = 3, x - y = 1 by graphing?

OpenStudy (tkhunny):

You could try graphing the two linear equations.

OpenStudy (anonymous):

And... how do I do that.

OpenStudy (tkhunny):

Define some coordinate axes. Have youheard of the Slope-Intercept Form of a line?

OpenStudy (anonymous):

Uhh. I don't believe so. is that y=mx+b?

OpenStudy (tkhunny):

Yes. y = -x + 3 y = x - 1 Let's see your best art work.

OpenStudy (anonymous):